OpenStudy (anonymous):

Simplify r^2 s^5 t^4/ r^2 s^2 t^5?

OpenStudy (anonymous):

When dividing with two same bases and different exponents, subtract \[\frac{ 2^5 }{ 2^2 } = \ 2^(5-2) = 2^3 = 8\]

OpenStudy (anonymous):

Likewise, \[\frac{ x^2 }{ x^2 } = 1\]

OpenStudy (anonymous):

Do that first, then we will handle the negative exponent you should get for t

OpenStudy (anonymous):

so, r would be eliminated, s^3, and t^-1?

OpenStudy (jhannybean):

Just group them:\[\large \frac{s^5}{s^2} \cdot \frac{r^2}{r^2} \cdot \frac{t^4}{t^5}\]And remember, \[\large \frac{x^m}{x^n} = x^{m-n}\]\[x^{-1}\iff \frac{1}{x}\]